RAMAN AND INFRARED SPECTRA, CONFORMATIONAL STABILITY, BARRIERS TO INTERNAL RO-TATION, AB INITIO CALCULATIONS AND r_{o} STRUCTURE FOR VINYL SILYL FLUORIDE

\underline{YASSER NASHED}, GAMIL A. GUIRGIS, M. A. QTAITAT AND J. R. DURIG, Department of Chemistry, University of Missouri-Kansas City, Kansas City, MO 64110-2499.

The Raman (3250 to 10 cm^{−1}) and infrared (3250 to 40 cm^{−1}) spectra of the gaseous and solid vinyl silyl fluoride (CH_{2}CHSiH_{2}F), have been recorded. From the far infrared spectrum of the gas, the fundamental asymmetric torsions for both the cis and gauche conformers have been observed at 102.34 and 86.56 cm^{−1}, respectively, with each having several excited state transitions falling to lower frequencies. From these transitions, the potential function to internal rotation has been determined with the following values: V_{1}, V_{2}, V_{3}, V_{4} and V_{6}, with the gauche conformer thermodynamically preferred by 82±20 cm^{−1}. The cis to gauche, gauche to gauche, and gauche to cis barriers are 702 (8.40 kJ/mol), 588 (7.04 kJ/mol) and 620 cm^{−1} (7.42 kJ/mol), respectively. Variable temperature (-105 to -150^{∘}C) studies of the infrared spectra of the sample dissolved in liquid krypton have been carried out. From these data, the enthalpy difference has been determined to be 76 cm^{−1} (0.91 kJ/mol). Additionally, ab initio calculations have been carried out utilizing the 6-31G(d) basis sets to obtain the conformational stability, barriers to internal rotation and optimized structural parameters.